The Equations are:
2500x+400y=19050
1200x+500y=1900
I need to find both x and y, but it also needs to be solved using the Elimination/Combination method. Everytime I try it, I get a different answer, and some of the numbers look way out of range when I do it.

Answer by checkley71(8403)

250X+40Y=1905
12X+5Y=19
250X+40Y=1905
-96X-40Y=-152 (MULTIPLIED BY -8)
---------------
154X =1753 OR X=1753/154 SUBSTITUTING X=1753/154 IN FORMULA #2 WE GET
12(1753/154)+5Y=19 OR 21036/154+5Y=19 OR 5Y=19-21036/154 OR
5Y=(2926-21036)/154 OR 5Y=-18110/154 OR Y=(-18110/154)/5 OR Y=(-18110/154)1/5
OR Y=-18110/770 OR Y=-1811/77
PROOF 12(1753/154)+5(-1811/77)=19 OR 21036/154-9055/77=19 OR
21O36/154-18110/154=19 OR 2926/154=19 OR 19=19.